# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ IV.2.2 Electron-doping-induced states
where $\mu_{+}$ denotes the $\omega$ value at the bottom of the upper band:
@@
In electron-doped Mott insulators of the HMs, because the upper band in the momentum regime outside the Fermi sea has essentially no electrons, the electronic states emerge in the momentum regime primarily inside the Fermi sea in the electron-removal spectrum.
In the 1D and 2D HMs, the emergence of electronic modes by electron doping in accordance with the above selection rules is confirmed as follows: by representing the excitation energy of the spin mode as $e^{\cal T}_{\bm{k}}$ [Figs. 1(b), 1(e), 2(b), and 2(c)] and noting that the Fermi momentum of the electron-doped system is ${\bm{k}}^{+}_{\rm F}=\frac{\bm{\pi}}{2}$ in the small-doping limit [Figs. 1(a), 1(d), and 2(a)], according to Eq. (48), an electronic mode can emerge from the bottom of the upper band ($\mu_{+}$), exhibiting the dispersion relation of
@@ IV.3 Electronic excited states from a perturbed state in the effective theory
where $i_{1}$ and $i_{2}$ represent the sites in the unit cell $i$.
@@ IV.4.1 Conventional bands in doped systems
IV.4.2 Ingap doping-induced states originating from spin-excited states
@@
The dispersion relations of the ingap hole-doping-induced mode [Eq. (65)] are shown by the solid red curves in Figs. 8(g), 8(j), 8(s), 8(v), 9(g), and 9(j), and those of the ingap electron-doping-induced mode [Eq. (66)] are shown by the solid magenta curves in Figs. 8(h), 8(k), 8(t), 8(w), 9(h), and 9(k), which are consistent with the numerical results [Figs. 8(a), 8(b), 8(d), 8(e), 8(m), 8(n), 8(p), 8(q), 9(a), 9(b), 9(d), and 9(e)].
@@
Figure 8: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/(2N_{\rm u})$ in the ladder HM for $U/t=8$ and $t_{\perp}/t=4$ [(a)–(c), (g)–(i)], bilayer HM for $U/t=16$ and $t_{\perp}/t=8$ [(d)–(f), (j)–(l)], 1D PAM for $U/t=14$ and $t_{\rm K}/t=4$ [(m)–(o), (s)–(u)], and 2D PAM for $U/t=28$ and $t_{\rm K}/t=8$ [(p)–(r), (v)–(x)]. (a)–(f), (m)–(r) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)–(c)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)–(f)], $A_{X}(k,\omega)t$ [(m)–(o)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(p)–(r)] for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=-\eta$] [(a), (d), (m), (p)], electron-doped ground state at $\delta=-0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=\eta$] [(b), (e), (n), (q)], and lowest-energy state of $\eta/(2N_{\rm u})=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c), (f), (o), (r)]. Gaussian broadening is used with standard deviations of $0.16t$ [(a)–(c), (m)–(o)] and $0.32t$ [(d)–(f), (p)–(r)]. (g)–(l), (s)–(x) Dispersion relations from the lowest-energy state of $\eta/(2N_{\rm u})\rightarrow+0$ for $\delta\rightarrow+0$ [(g), (j), (s), (v)], $\delta\rightarrow-0$ [(h), (k), (t), (w)], and $\delta=0$ [(i), (l), (u), (x)] in the effective theory. The dispersion relations of the emergent modes are $\omega=e^{{\cal T}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ (solid red curves), $\omega=e^{{\cal C}^{0}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ for $\omega>0$ (dotted red curves), and $\omega=-e^{{\cal C}^{-1}}_{-{\bm{k}}_{\parallel}-{\bm{\pi}}}-\mu_{-}$ for $\omega<0$ (dotted magenta curves) in (g), (i), (j), (l), (s), (u), (v), and (x); $\omega=-e^{{\cal T}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ (solid magenta curves), $\omega=-e^{{\cal C}^{0}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ for $\omega<0$ (dotted red curves), and $\omega=e^{{\cal C}^{1}}_{{\bm{k}}_{\parallel}}-\mu_{+}$ for $\omega>0$ (dotted magenta curves) in (h), (i), (k), (l), (t), (u), (w), and (x), which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i), (j)–(l), (s)–(u), and (v)–(x) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively, with $\mu$ adjusted such that $\mu_{-}=0$ [(g), (j), (s), (v)], $\mu_{+}=0$ [(h), (k), (t), (w)], and $\mu=0$ [(i), (l), (u), (x)]. The solid green lines indicate $\omega=0$.
@@
Figure 9: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/(2N_{\rm u})$ in the 1D KLM for $J_{\rm K}/t=4$ [(a)–(c), (g)–(i)] and 2D KLM for $J_{\rm K}/t=8$ [(d)–(f), (j)–(l)]. (a)–(f) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=-\eta$] [(a), (d)], electron-doped ground state at $\delta=-0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=\eta$] [(b), (e)], and lowest-energy state of $\eta/(2N_{\rm u})=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c), (f)]. Gaussian broadening is used with standard deviations of $0.12t$ [(a)–(c)] and $0.24t$ [(d)–(f)]. (g)–(l) Dispersion relations from the lowest-energy state of $\eta/(2N_{\rm u})\rightarrow+0$ for $\delta\rightarrow+0$ [(g), (j)], $\delta\rightarrow-0$ [(h), (k)], and $\delta=0$ [(i), (l)] in the effective theory. The dispersion relations of the emergent modes are $\omega=e^{{\cal T}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ (solid red curves) [(g), (i), (j), (l)] and $\omega=-e^{{\cal T}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ (solid magenta curves) [(h), (i), (k), (l)], which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i) and (j)–(l) are the same as those in Figs. 3(i) and 3(l), respectively, with $\mu$ adjusted such that $\mu_{-}=0$ [(g), (j)], $\mu_{+}=0$ [(h), (k)], and $\mu=0$ [(i), (l)]. The solid green lines indicate $\omega=0$.
@@ IV.5 Spectral weights and dispersion relations of doping-induced states in the small-doping regime
In the $n$-hole-doped ($n$-electron-doped) system for $n\ll N_{\rm u}$, because there are essentially $n$ doped unit cells $|X\rangle_{i}=|{\cal R}_{\sigma^{\prime}}\rangle_{i}$ ($|{\cal A}_{\sigma^{\prime}}\rangle_{i}$) in the second terms on the right-hand sides of Eqs. (IV.3) and (IV.3), the probability of adding (removing) an electron on the doped unit cells is essentially $n$ times as large as that for the 1-hole-doped (1-electron-doped) system; the spectral weights of the doping-induced states essentially become $n$ times as large as those of the 1-hole-doped (1-electron-doped) system. This means that the spectral weights of the doping-induced states increase proportionally to the doping concentration $|\delta|$ in the small-$|\delta|$ regime and that the doping-induced modes and changes in the band structure can be observed, provided that a macroscopic number $\mathcal{O}(N_{\rm u})$ of holes or electrons can be introduced via doping KohnoDIS ; KohnoKLM ; KohnoTinduced ; KohnoRPP ; Eskes ; DagottoDOS .
@@ IV.6 Emergent states corresponding to doping-induced states in undoped charge-perturbed systems
using the spectral functions of the $2n$-hole-doped and $2n$-electron-doped ground states for $\mu=0$ [Appendix A.2; Eq. (A.2)]. This state ($|\eta=n,\eta^{z}=0\rangle$) is the lowest-energy state of $\eta=n$ at half filling ($\eta^{z}=0$), because the doped ground states at $N_{\rm h}=\pm 2n$ are the lowest-energy states of $\eta=n$ and $\eta^{z}=\mp n$, provided that phase separation for the electron density does not occur at $N_{\rm h}=\pm 2n$ for $H=S^{z}=0$.
@@
The dispersion relations of electronic excitations from the lowest-energy charge-perturbed state in the small $\eta$-density limit at half filling obtained via Eq. (72) using the Bethe ansatz and effective theory are shown in Figs. 6(f), 8(i), 8(l), 8(u), 8(x), 9(i), and 9(l). The spectral functions of the state corresponding to the lowest-energy charge-perturbed state of small $\eta$-density at half filling obtained via Eq. (72) using numerical calculations are shown in Figs. 6(c), 7(c), 8(c), 8(f), 8(o), 8(r), 9(c), and 9(f).
@@ IV.6.2 Remarks on nonzero-$\eta$ states at half filling
In the case of photon (light) radiation, because a particle-hole pair can be excited by a photon, a macroscopic number $\mathcal{O}(N_{\rm u})$ of photons that excite the particle-hole pairs (excluding reflected and transmitted photons which are irrelevant to the excitations) should be injected every lifetime (relaxation time) of the particle-hole excited states to keep the number of excited particle-hole pairs macroscopic, as in a steady state, during the measurement.
@@ IV.7.1 Selection rules for electronic excitation from a state perturbed by charge fluctuation
Electronic modes can also be induced by charge fluctuation. As a charge-perturbed state, we consider ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ at half filling for $\lambda=c$ and $f$ [Eq. (15)]. The charge-perturbed state ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of $\bm{q}$}, $N_{\rm h}=0$, and $S^{z}=0$ for the unperturbed ground state $|{\rm GS}\rangle$ with the momentum of $\bm{0}$}, $N_{\rm h}=0$, and $S^{z}=0$. The excitation energy of ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ from $|{\rm GS}\rangle$ is denoted as $e^{{\cal C}^{0}}_{\bm{q}}$.
The electron-addition ($N_{\rm h}=-1$) and electron-removal ($N_{\rm h}=1$) excited states with a momentum $\bm{k}$} in the conventional bands from $|{\rm GS}\rangle$ are represented as $|{\tilde{\cal A}}\rangle_{\bm{k}}$ and $|{\tilde{\cal R}}\rangle_{\bm{k}}$; their excitation energies from $|{\rm GS}\rangle$ are denoted as $\varepsilon^{\tilde{\cal A}}_{\bm{k}}$ and $\varepsilon^{\tilde{\cal R}}_{\bm{k}}$, respectively. The dispersion relations of the electronic modes from $|{\rm GS}\rangle$ are $\omega=\varepsilon^{\tilde{\cal A}}_{\bm{k}}$ for $\omega>0$ and $\omega=-\varepsilon^{\tilde{\cal R}}_{-{\bm{k}}}$ for $\omega<0$ [Figs. 1(a), 1(d), 2(a), and 3].
The state obtained by adding an electron with a momentum $\bm{k}$} to ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=-1$. This state can overlap with $|{\tilde{\cal A}}\rangle_{{\bm{k}}+{\bm{q}}}$ and emerge in the electron-addition spectrum, exhibiting
@@
Similarly, the state obtained by removing an electron with a momentum $\bm{k}$} from ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=1$. This state can overlap with $|{\tilde{\cal R}}\rangle_{-{\bm{k}}+{\bm{q}}}$ and emerge in the electron-removal spectrum, exhibiting

### v2

@@ IV.2.2 Electron-doping-induced states
where $\mu_{+}$ denotes the $\omega$ value at the bottom of the upper band,
@@
In electron-doped Mott insulators of the HM, because the upper band in the momentum regime outside the Fermi sea has essentially no electrons, the electronic states emerge in the momentum regime primarily inside the Fermi sea in the electron-removal spectrum.
In the 1D and 2D HMs, the emergence of electronic modes caused by electron doping in line with the above selection rules is confirmed as follows: by representing the excitation energy of the spin mode as $e^{\cal T}_{\bm{k}}$ [Figs. 1(b), 1(e), 2(b), and 2(c)] and noting that the Fermi momentum of the electron-doped system is ${\bm{k}}^{+}_{\rm F}=\frac{\bm{\pi}}{2}$ in the small-doping limit [Figs. 1(a), 1(d), and 2(a)], according to Eq. (50), an electronic mode can emerge from the bottom of the upper band ($\mu_{+}$), exhibiting the dispersion relation of
@@ IV.3 Electronic excited states from a perturbed state in the effective theory
where $i_{m}$ represents the site within the unit cell $i$ on the $m$th leg in the ladder HM and on the $m$th plane in the bilayer HM.
@@ IV.4.1 Conventional bands in doped systems
IV.4.2 In-gap doping-induced states originating from spin-excited states
@@
The dispersion relations of the in-gap hole-doping-induced mode [Eq. (67)] are shown by the solid red curves in Figs. 8(g), 8(j), 8(s), 8(v), 9(g), and 9(j), and those of the in-gap electron-doping-induced mode [Eq. (68)] are shown by the solid magenta curves in Figs. 8(h), 8(k), 8(t), 8(w), 9(h), and 9(k), which are consistent with the numerical results [Figs. 8(a), 8(b), 8(d), 8(e), 8(m), 8(n), 8(p), 8(q), 9(a), 9(b), 9(d), and 9(e)].
@@
Figure 8: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/(2N_{\rm u})$ in the ladder HM [(a)–(c), (g)–(i)], bilayer HM [(d)–(f), (j)–(l)], 1D PAM [(m)–(o), (s)–(u)], and 2D PAM [(p)–(r), (v)–(x)]. (a)–(f), (m)–(r) $A_{X}(k_{\parallel},0,\omega)t+A_{X}(k_{\parallel},\pi,\omega)t$ [(a)–(c)], ${\bar{A}}_{X}({\bm{k}}_{\parallel},0,\omega)t+{\bar{A}}_{X}({\bm{k}}_{\parallel},\pi,\omega)t$ [(d)–(f)], $A_{X}(k,\omega)t$ [(m)–(o)], and ${\bar{A}}_{X}({\bm{k}},\omega)t$ [(p)–(r)] for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=-\eta$] [(a), (d), (m), (p)], electron-doped ground state at $\delta=-0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=\eta$] [(b), (e), (n), (q)], and lowest-energy state of $\eta/(2N_{\rm u})=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c), (f), (o), (r)]. (g)–(l), (s)–(x) Dispersion relations from the lowest-energy state of $\eta/(2N_{\rm u})\rightarrow+0$ for $\delta\rightarrow+0$ [(g), (j), (s), (v)], $\delta\rightarrow-0$ [(h), (k), (t), (w)], and $\delta=0$ [(i), (l), (u), (x)] in the effective theory. The dispersion relations of the emergent modes are $\omega=e^{{\cal T}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ (solid red curves), $\omega=e^{{\cal C}^{0}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ for $\omega>0$ (dotted red curves), and $\omega=-e^{{\cal C}^{-1}}_{-{\bm{k}}_{\parallel}-{\bm{\pi}}}-\mu_{-}$ for $\omega<0$ (dotted magenta curves) in (g), (i), (j), (l), (s), (u), (v), and (x); $\omega=-e^{{\cal T}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ (solid magenta curves), $\omega=-e^{{\cal C}^{0}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ for $\omega<0$ (dotted red curves), and $\omega=e^{{\cal C}^{1}}_{{\bm{k}}_{\parallel}}-\mu_{+}$ for $\omega>0$ (dotted magenta curves) in (h), (i), (k), (l), (t), (u), (w), and (x), which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i), (j)–(l), (s)–(u), and (v)–(x) are the same as those in Figs. 3(g), 3(j), 3(h), and 3(k), respectively, with $\mu$ adjusted such that $\mu_{-}=0$ [(g), (j), (s), (v)], $\mu_{+}=0$ [(h), (k), (t), (w)], and $\mu=0$ [(i), (l), (u), (x)]. The solid green lines indicate $\omega=0$.
@@
Figure 9: Electronic excitation from the lowest-energy charge-perturbed states of small $\eta/(2N_{\rm u})$ in the 1D KLM [(a)–(c), (g)–(i)] and 2D KLM [(d)–(f), (j)–(l)]. (a)–(f) $A_{X}({\bm{k}},\omega)t$ for $|X\rangle$ of the state corresponding to the hole-doped ground state at $\delta=0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=-\eta$] [(a), (d)], electron-doped ground state at $\delta=-0.05$ $\eta/(2N_{\rm u})=0.025$ and $\eta^{z}=\eta$] [(b), (e)], and lowest-energy state of $\eta/(2N_{\rm u})=0.025$ at $\delta=0$ ($\eta^{z}=0$) [(c), (f)]. (g)–(l) Dispersion relations from the lowest-energy state of $\eta/(2N_{\rm u})\rightarrow+0$ for $\delta\rightarrow+0$ [(g), (j)], $\delta\rightarrow-0$ [(h), (k)], and $\delta=0$ [(i), (l)] in the effective theory. The dispersion relations of the emergent modes are $\omega=e^{{\cal T}}_{{\bm{k}}_{\parallel}-{\bm{\pi}}}+\mu_{-}$ (solid red curves) [(g), (i), (j), (l)] and $\omega=-e^{{\cal T}}_{-{\bm{k}}_{\parallel}}+\mu_{+}$ (solid magenta curves) [(h), (i), (k), (l)], which are shown in the relevant momentum regimes for comparison with the numerical results. The dispersion relations of the other modes and continua in (g)–(i) and (j)–(l) are the same as those in Figs. 3(i) and 3(l), respectively, with $\mu$ adjusted such that $\mu_{-}=0$ [(g), (j)], $\mu_{+}=0$ [(h), (k)], and $\mu=0$ [(i), (l)]. The solid green lines indicate $\omega=0$.
@@ IV.5 Spectral weights and dispersion relations of doping-induced states in the small-doping regime
In the $n$-hole-doped ($n$-electron-doped) system for $n\ll N_{\rm u}$, because there are essentially $n$ doped unit cells $|X\rangle_{i}=|{\cal R}_{\sigma^{\prime}}\rangle_{i}$ ($|{\cal A}_{\sigma^{\prime}}\rangle_{i}$) in the second terms on the right-hand sides of Eqs. (IV.3) and (IV.3), the probability of adding (removing) an electron on the doped unit cells is essentially $n$ times as large as that for the one-hole-doped (one-electron-doped) system; the spectral weights of the doping-induced states essentially become $n$ times as large as those of the one-hole-doped (one-electron-doped) system. This means that the spectral weights of the doping-induced states increase proportionally to the doping concentration $|\delta|$ in the small-$|\delta|$ regime and that the doping-induced modes and changes in the band structure can be observed, provided that a macroscopic number $\mathcal{O}(N_{\rm u})$ of holes or electrons can be introduced via doping KohnoDIS ; KohnoKLM ; KohnoTinduced ; KohnoRPP ; Eskes ; DagottoDOS .
@@ IV.6 Emergent states corresponding to doping-induced states in undoped charge-perturbed systems
using the spectral functions of the $2n$-hole-doped and $2n$-electron-doped ground states for $\mu=0$ [Appendix A.2; Eq. (A.2)]. This state ($|\eta=n,\eta^{z}=0\rangle$) is the lowest-energy state of $\eta=n$ at half filling ($\eta^{z}=0$), because the doped ground states at $N_{\rm h}=\pm 2n$ are the lowest-energy states of $\eta=n$ and $\eta^{z}=\mp n$, provided that phase separation in the electron density does not occur at $N_{\rm h}=\pm 2n$ for $H=S^{z}=0$.
@@
The dispersion relations of electronic excitations from the lowest-energy charge-perturbed state in the small $\eta$-density limit at half filling obtained via Eq. (74) using the Bethe ansatz and the effective theory are shown in Figs. 6(f), 8(i), 8(l), 8(u), 8(x), 9(i), and 9(l). The spectral functions of the state corresponding to the lowest-energy charge-perturbed state of small $\eta$-density at half filling obtained via Eq. (74) using numerical calculations are shown in Figs. 6(c), 7(c), 8(c), 8(f), 8(o), 8(r), 9(c), and 9(f).
@@ IV.6.2 Remarks on nonzero-$\eta$ states at half filling
In the case of photon (light) radiation, because a particle-hole pair can be excited by a photon, a macroscopic number $\mathcal{O}(N_{\rm u})$ of photons that excite the particle-hole pairs (excluding reflected and transmitted photons that are irrelevant to the excitations) should be injected every lifetime (relaxation time) of the particle-hole excited states to keep the number of excited particle-hole pairs macroscopic, as in a steady state, during the measurement.
@@ IV.7.1 Selection rules for electronic excitation from a state perturbed by charge fluctuation
Electronic modes can also be induced by charge fluctuation. As a charge-perturbed state, we consider ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ at half filling for $\lambda=c$ and $f$ [Eq. (17)]. The charge-perturbed state ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of $\bm{q}$}, $N_{\rm h}=0$, and $S^{z}=0$ for the unperturbed ground state $|{\rm GS}\rangle$ with the momentum of $\bm{0}$}, $N_{\rm h}=0$, and $S^{z}=0$. The excitation energy of each eigenstate in ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ from $|{\rm GS}\rangle$ is denoted by $e^{{\cal C}^{0}}_{\bm{q}}$.
The electron-addition ($N_{\rm h}=-1$) and electron-removal ($N_{\rm h}=1$) excited states with a momentum $\bm{k}$} in the conventional bands from $|{\rm GS}\rangle$ are represented by $|{\tilde{\cal A}}\rangle_{\bm{k}}$ and $|{\tilde{\cal R}}\rangle_{\bm{k}}$; their excitation energies from $|{\rm GS}\rangle$ are denoted by $\varepsilon^{\tilde{\cal A}}_{\bm{k}}$ and $\varepsilon^{\tilde{\cal R}}_{\bm{k}}$, respectively. The dispersion relations of the electronic modes from $|{\rm GS}\rangle$ are $\omega=\varepsilon^{\tilde{\cal A}}_{\bm{k}}$ for $\omega>0$ and $\omega=-\varepsilon^{\tilde{\cal R}}_{-{\bm{k}}}$ for $\omega<0$ [Figs. 1(a), 1(d), 2(a), and 3].
The state obtained by adding an electron with a momentum $\bm{k}$} to ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=-1$. Since this state can overlap with $|{\tilde{\cal A}}\rangle_{{\bm{k}}+{\bm{q}}}$, electronic states can emerge in the electron-addition spectrum, exhibiting
@@
Similarly, the state obtained by removing an electron with a momentum $\bm{k}$} from ${\tilde{n}}^{\lambda}_{\bm{q}}|{\rm GS}\rangle$ has the momentum of ${\bm{k}}+{\bm{q}}$ and $N_{\rm h}=1$. Since this state can overlap with $|{\tilde{\cal R}}\rangle_{-{\bm{k}}+{\bm{q}}}$, electronic states can emerge in the electron-removal spectrum, exhibiting